OpenStudy (brekathleen):

please help solve for x : 3|2x-2|+6=18

OpenStudy (mathmale):

this is an "absolute value" problem. |x-1| = 5 can be decomposed (broken down) into two separate equations: 1) +(X-1) = 5 2) - (X-1) = 5 Try solving these two equations for x. Check your answers. Then go back to the problem you've posted and apply your knowledge to that. Hint: Div. both sides of the given equation by 3 first.

OpenStudy (mathstudent55):

Can you first simplify your given equation to have only an absolute value equaling a number? To do that, you need to get rid of the 6 and the 3 on the left side. First, subtract 6 from both sides. Then divide both sides by 3. Can you do that and show what you got?

jhonyy9 (jhonyy9):

so the first step substract from both sides 6 and so will get 3I2x-2I +6-6=18-6 3I2x-2I=12 so now divide both sides by 3 and will get I2x-2I=4 - so and now you can using what you know about I2x-2I= ? so these two cases